# VTarbel: Targeted Label Attack with Minimal Knowledge on Detector-enhanced Vertical Federated Learning

### v1

@@ 1. Introduction
Vertical Federated Learning (VFL) is a decentralized machine learning (ML) paradigm that enables collaborative model training among multiple participants, each holding distinct, non-overlapping feature sets, without sharing their raw data  (Yang et al., 2019; Liu et al., 2022a; Li et al., 2023; Yang et al., 2023; Wei et al., 2022; Khan et al., 2022; Jia et al., 2024). Typically, one participant, referred to as the active party, possesses the ground-truth labels, while the others, referred to as passive parties, contribute only raw features. VFL has gained traction in real-world applications such as financial risk assessment  (Chen et al., 2021), healthcare diagnostics  (Yang et al., 2019), and personalized advertising  (Li et al., 2022b).
@@
Despite growing attention to privacy breaches in VFL, its robustness against adversarial perturbations remain underexplored. A particularly critical but overlooked threat is targeted label attacks, in which a passive party (the attacker) deliberately crafts malicious input samples to mislead the VFL model into producing attacker-specified misclassifications. For instance, an adversary in a healthcare VFL system could manipulate diagnostic outputs to misguide treatments. Such attacks generally fall into two categories: (1) Training-phase backdoor attacks (Naseri et al., 2024; Bai et al., 2023; Liu et al., 2024), where attackers poison training data to implant hidden triggers, once activated during inference, cause targeted misclassifications. (2) Inference-phase adversarial attacks (Gu and Bai, 2023; Qiu et al., 2024; Pang et al., 2023; He et al., 2023; Liu et al., 2022b), where attackers directly perturb test inputs to manipulate model predictions without modifying the training process (as illustrated in Figure 1). While both types pose security risks, backdoor attacks often degrade overall model performance due to the presence of poisoned samples, potentially alerting the active party (the defender) and lead to the early termination of the collaboration. In contrast, adversarial attacks operate stealthily without altering training data, which significantly increases the difficulty of detection. Given their practicality and stealthiness, this work focuses on inference-phase targeted label attacks (hereafter refer to as targeted label attacks for brevity).
@@ 2.1. Detector-enhanced VFL Inference System
A detector-enhanced VFL inference system, as depicted in Figure 1, comprises $K$ participants $P_{1},P_{2},\ldots,P_{K}$. We assume $P_{K}$ serves as the only active party, while the remaining $K-1$ participants act as passive parties. Each participant $P_{k}$ ($k\in[K]$, where $[K]=\{1,2,\ldots,K\}$) possesses a private test dataset $D_{k}=\{X_{k}^{i}\}_{i=1}^{N}$, where $X_{k}^{i}$ denotes the feature vector of the $i$-th test sample for $P_{k}$, and $N$ represents the total number of test samples.
@@
In a conventional VFL system (without defenses), $E$ is directly input to the top model $g(\theta_{T})$ to generate the final prediction $\hat{y}=g(E;\theta_{T})$, where $\hat{y}\in[C]$ and $C$ is the number of classes. This prediction is then returned to all passive parties. However, practical deployments face risks from malicious passive parties (e.g., $P_{1}$ in Figure 1) that may submit adversarial embeddings (e.g., $e_{1}$) to manipulate predictions via targeted label attacks. To mitigate this, the detector-enhanced system introduces an anomaly detector $\phi(\cdot)$ at $P_{K}$. After computing $\hat{y}$, the system evaluates $\phi(E,\hat{y})$. If $E$ is flagged as anomalous, the detector triggers a rejection by returning the special label REJ to all participants; otherwise, $\hat{y}$ is returned as usual. For consistency, we treat REJ as an additional class, extending the total number of classes to $C+1$.
@@ 4.1. Observations
Our experiments are conducted within a two-party VFL inference system, where the passive party is compromised by the attacker and the active party acts as the defender, employing DeepAE as the anomaly detector. We evaluate three different bottom model architectures: MLP3, VGG13 (Simonyan and Zisserman, 2014), and ResNet18 (He et al., 2016), across two datasets: TabFMNIST (Xiao et al., 2017) and CIFAR10 (Krizhevsky et al., 2014). Without loss of generality, we set the targeted label $t^{\star}=0$ for both datasets. The experimental results are presented in Figure 4, which reveals two key observations.
Observation 1: High ASR without detectors. As shown in Figure 4, for all three attack methods, the ASR is significantly higher in the absence of anomaly detectors compared to the “Ground-Truth” case. For instance, in the configuration (MLP3, TabFMNIST), the ASR increases from 10.00% in the “Ground-Truth” case to 29.39%, 83.68%, and 97.81% for LR-BA, HijackVFL, and ADI, respectively. These results indicate that the feature vectors, maliciously optimized against the attacker’s local auxiliary model, effectively transfer to the global VFL model.
Observation 2: Significant drop in attack performance with detectors. Figure 4 also shows a drastic decline in ASR when detectors are deployed by the defender. In some cases, the ASR even falls to zero, which is even lower than the baseline “Ground-Truth” performance. For instance, in the (ResNet18, CIFAR10) configuration, the ASR for LR-BA decreases from 91.40% (without detector) to 31.84% (with detector). This drop is due to the anomaly detector recognizing many malicious feature embeddings as anomalies, causing the global VFL model to predict the special rejection label REJ. In other configurations, such as (VGG13, CIFAR10), the ASRs for LR-BA and HijackVFL with detectors are 0.04% and 0, respectively—significantly lower than the 10.0% achieved in the “Ground-Truth” case. This highlights the ineffectiveness of these attacks in practical VFL scenarios where anomaly detectors are deployed.
@@
Figure 4. Impact of anomaly detector on ASR in VFL inference system. The “Ground-Truth” represents the proportion of the targeted label in the original test set, serving as a baseline for comparison.
@@ 4.2. Main Idea of VTarbel
In summary, we propose a novel two-stage attack framework, VTarbel, which will be described in detail in the next section. This framework generalizes previous methods, where the “Ground-Truth” case and previous attacks illustrated in Figure  4 are special cases of our framework. Specifically, the “Ground-Truth” corresponds to using only the first stage without performing attacks in the second stage, while previous attacks correspond to cases where only the second stage is applied without the first stage. In Section  6.3, we demonstrate that both first-stage-only and second-stage-only approaches are ineffective, and only our generalized two-stage framework achieves the highest ASR.
@@ 5.1. Two-stage Attack Formulation
From the above analysis, we identify a fundamental trade-off in the design of the two-stage attack framework. From Eq.(9), it is evident that the total number of test samples predicted as targeted label is the summation of two terms: the number of samples predicted as targeted label in preparation stage and the number predicted in attack stage. If the preparation stage is too short, although a large number of test samples remain for the attack stage, the relatively low ASR $s_{2}$ will result in a small second term in Eq.(9), leading to an ineffective attack. All previous attack methods fall into this category, corresponding to the special edge case where the preparation stage is empty, i.e., $\mathcal{Q}=\emptyset$. On the other hand, if the preparation stage is too long, the ASR $s_{2}$ will be high, but the number of test samples remaining for the attack stage will be small. This also causes the second term in Eq.(9) to be small, resulting in an ineffective attack. The “Ground-Truth” case in Figure  4 falls into this category, corresponding to the special edge case where $\mathcal{Q}=D_{adv}$.
The experimental results in Figure  6 confirm our analysis. The overall ASR $s$ initially improves as $\rho$ increases from 1% to 5%. However, when $\rho$ becomes relatively large (10% to 90%), there is a corresponding decrease in $s$. To effectively address this trade-off, our design goal is to construct a preparation stage that maximizes $s_{2}$ while keeping its length minimal, thereby preserving as many test samples as possible for the attack stage.
@@ 6.4. Optimality Verification of Preparation Set
We observe that, in most cases, the preparation set selected by VTarbel achieves the optimal ASR. For instance, under the (ResNet18, CIFAR10)configuration, VTarbel achieves the highest ASR of 90.28%, while the early stopping ($-15\%$) and late stopping ($+15\%$) variants only achieve ASRs of 81.97% and 83.28%, respectively, both of which are substantially lower than VTarbel’s result. The reason behind this performance difference is as follows: For late stopping, the optimal $|\mathcal{Q}^{*}|$ size provides enough test samples for the surrogate model and estimated detector to be sufficiently trained. Adding more samples in the preparation stage reduces the number of samples predicted as the targeted label during the attack stage, thus lowering the ASR. In contrast, for early stopping, fewer test samples than VTarbel’s optimal preparation set slightly degrade the performance of the estimated detector, leading to a reduced number of samples predicted as targeted label in the attack stage, which also results in a lower ASR compared to VTarbel.
@@ 6.5. Detector-Agnostic Attack
Figure 9. The impact of the attacker’s feature fraction in a two-participant VFL inference system, where the features allocated to each participant are randomly selected.
@@ 6.6. Scalability Evaluation
In this section, we evaluate VTarbel’s attack scalability in terms of the attacker’s feature fraction and the total number of participants in the VFL inference system. Our experiments are conducted using the TabMNIST and TabFMNIST datasets under two scenarios: the first scenario corresponds to a VFL inference system with two participants, where the attacker’s feature fraction (i.e., the fraction of the attacker’s features to the total number of features) ranges from 90% to 10% (Figure 9). The second scenario corresponds to a VFL inference system with varying numbers of participants, ranging from 2 to 10, where the total features are randomly and evenly allocated to each participant (Figure 11).
From both Figure 9 and Figure 11, several similar trends can be observed. First, the baseline ASR remains stable and is unaffected by the attacker’s feature fraction or the number of participants. This is because the baseline is only related to the global VFL model’s accuracy. Next, as the number of features allocated to the attacker decreases (i.e., the decrease in the attacker’s feature fraction in Figure 9 and the increase in the total number of participants in Figure 11), both the accuracy of the locally trained surrogate model and the F1 score of the estimated detector gradually decrease. The reason for this is that with fewer features, it becomes much more challenging for the attacker to leverage these limited features to train an accurate surrogate model or detector that can precisely estimate the global VFL model or the defender’s detector. Consequently, there is a corresponding gradual decrease in VTarbel’s ASR, which is influenced not only by the less accurate surrogate model and detector but also by the benign features from other participants that can neutralize the malicious properties of the attacker’s feature embedding. However, we note that practical VFL systems typically involve only two participants, where the passive party (i.e., the attacker) serves as the feature provider and often possesses more features than the active party (i.e., the defender) (Chen et al., 2021). In such practical settings, VTarbel demonstrates strong attack performance. For instance, as shown in Figure  9, when the attacker holds 60% of the features, the ASR reaches 91.97% on TabMNIST and 90.73% on TabFMNIST, posing a significant threat to the security of practical VFL inference systems.
Figure 11. The impact of the number of participants, where the total features are randomly and evenly split among participants.
Figure 13. Impact of an imbalance binary dataset, where the negative class is the targeted label.
@@ 6.7.1. Imbalanced Dataset
We investigate the impact of an imbalanced dataset on VTarbel’s attack performance. We construct a binary test set containing 1,000 samples by randomly selecting test samples from class 0 (treated as the negative class) and class 1 (treated as the positive class) from the CIFAR10 dataset, and evaluate it using the VGG13 and ResNet18 models. As shown in Figure 13, the imbalance ratio on the x-axis is defined as the number of negative samples divided by the number of positive samples, where the targeted label is the negative class. For VTarbel, we observe that even as the imbalance ratio increases from 1:1 to 1:8, the ASR remains stable and comparable to that of the balanced dataset. For example, when the imbalance ratio is 1:8, the ASR for VGG13 and ResNet18 is 0.916 and 0.919, respectively, both of which are close to the ASRs of 0.942 and 0.934 observed when the imbalance ratio is 1:1. This demonstrates that VTarbel is robust to class imbalance within the test set. Furthermore, when compared to the variant Random-Preparation (w/o clustering), we find that when the class imbalance becomes more severe (e.g., the imbalance ratio is greater than 1:5), the ASR for Random-Preparation (w/o clustering) is close to zero and even lower than that of the baseline. In contrast, VTarbel consistently achieves a considerably higher ASR than Random-Preparation (w/o clustering). This further validates that the clustering step in the preparation stage is crucial, as it ensures that the preparation set remains balanced, which facilitates the training of both the surrogate model and the estimated detector, even when the original test set is imbalanced.
Figure 16. Impact of anomaly detection threshold under the (ResNet18, CIFAR10, DeepAE) configuration.
@@ 6.7.2. Anomaly Detection Threshold
As shown in Figure 16, we plot the overall ASR of VTarbel, along with the ASR of the attack stage (since the detection thresholds have no effect on the ASR of the preparation stage) and the overall anomaly ratio identified by the defender’s detector. In Figure LABEL:fig:active_percentile, as we gradually increase the strength of the defender’s anomaly detector (e.g., with detection threshold percentiles lower than 50), there is a gradual increase in the anomaly ratio. This occurs because most of the attacker’s optimized feature embeddings are recognized as anomalies. As a result, the ASR of the attack stage gradually decreases, and the overall ASR of VTarbel also decreases simultaneously. In contrast, Figure LABEL:fig:passive_percentile shows that as the detection strength of the attacker’s estimated detector increases, the anomaly ratio remains stable and close to 0. The reason for this is that as the detection strength of the estimated detector increases, the optimization step in Algorithm 2 is terminated earlier with fewer optimization rounds, ensuring that the generated feature embedding stays within a region with high probability. Thus, the anomaly ratios consistently remain close to zero. At the same time, due to fewer optimization rounds, the malicious feature embeddings are less effective at being mapped to the targeted label, which explains the slight decrease in the ASR of the attack stage and the overall ASR of VTarbel.

### v2

@@ 1. Introduction
Vertical Federated Learning (VFL) is a decentralized machine learning (ML) paradigm that enables collaborative model training among multiple participants, each holding distinct, non-overlapping feature sets, without sharing their raw data  (Yang et al., 2019; Liu et al., 2022b; Li et al., 2023; Yang et al., 2023; Wei et al., 2022; Khan et al., 2022; Jia et al., 2024). Typically, one participant, referred to as the active party, possesses the ground-truth labels, while the others, referred to as passive parties, contribute only raw features. VFL has gained traction in real-world applications such as financial risk assessment  (Chen et al., 2021), healthcare diagnostics  (Yang et al., 2019), and personalized advertising  (Li et al., 2022b).
@@
Despite growing attention to privacy breaches in VFL, its robustness against adversarial perturbations remain underexplored. A particularly critical but overlooked threat is targeted label attacks, in which a passive party (the attacker) deliberately crafts malicious input samples to mislead the VFL model into producing attacker-specified misclassifications. For instance, an adversary in a healthcare VFL system could manipulate diagnostic outputs to misguide treatments. Such attacks generally fall into two categories: (1) Training-phase backdoor attacks (Naseri et al., 2024; Bai et al., 2023; Liu et al., 2024), where attackers poison training data to implant hidden triggers, once activated during inference, cause targeted misclassifications. (2) Inference-phase adversarial attacks (Gu and Bai, 2023; Qiu et al., 2024; Pang et al., 2023; He et al., 2023; Liu et al., 2022a), where attackers directly perturb test inputs to manipulate model predictions without modifying the training process (as illustrated in Figure 1). While both types pose security risks, backdoor attacks often degrade overall model performance due to the presence of poisoned samples, potentially alerting the active party (the defender) and lead to the early termination of the collaboration. In contrast, adversarial attacks operate stealthily without altering training data, which significantly increases the difficulty of detection. Given their practicality and stealthiness, this work focuses on inference-phase targeted label attacks (hereafter refer to as targeted label attacks for brevity).
@@ 2.1. Detector-enhanced VFL Inference System
A detector-enhanced VFL inference system, as depicted in Figure 1, comprises $K$ participants $P_{1},P_{2},\ldots,P_{K}$. We assume $P_{K}$ serves as the only active party, while the remaining K1K-1 participants act as passive parties. Each participant $P_{k}$ ($k\in[K]$, where $K]=\{1,2,\ldots,K\}$) possesses a private test dataset $D_{k}=\{X_{k}^{i}\}_{i=1}^{N}$, where $X_{k}^{i}$ denotes the feature vector of the $i$-th test sample for $P_{k}$, and $N$ represents the total number of test samples.
@@
[…]
@@ 4.1. Observations
Our experiments are conducted within a two-party VFL inference system, where the passive party is compromised by the attacker and the active party acts as the defender, employing DeepAE as the anomaly detector. We evaluate three different bottom model architectures: MLP3, VGG13 (Simonyan and Zisserman, 2014), and ResNet18 (He et al., 2016), across two datasets: TabFMNIST (Xiao et al., 2017) and CIFAR10 (Krizhevsky et al., 2014). Without loss of generality, we set the targeted label $t^{\star}=0$ for both datasets. The experimental results are presented in Figure 4(a), which reveals two key observations.
Observation 1: High ASR without detectors. As shown in Figure 4(a), for all three attack methods, the ASR is significantly higher in the absence of anomaly detectors compared to the “Ground-Truth” case. For instance, in the configuration (MLP3, TabFMNIST), the ASR increases from 10.00% in the “Ground-Truth” case to 29.39%, 83.68%, and 97.81% for LR-BA, HijackVFL, and ADI, respectively. These results indicate that the feature vectors, maliciously optimized against the attacker’s local auxiliary model, effectively transfer to the global VFL model.
Observation 2: Significant drop in attack performance with detectors. Figure 4(a) also shows a drastic decline in ASR when detectors are deployed by the defender. In some cases, the ASR even falls to zero, which is even lower than the baseline “Ground-Truth” performance. For instance, in the (ResNet18, CIFAR10) configuration, the ASR for LR-BA decreases from 91.40% (without detector) to 31.84% (with detector). This drop is due to the anomaly detector recognizing many malicious feature embeddings as anomalies, causing the global VFL model to predict the special rejection label REJ. In other configurations, such as (VGG13, CIFAR10), the ASRs for LR-BA and HijackVFL with detectors are 0.04% and 0, respectively—significantly lower than the 10.0% achieved in the “Ground-Truth” case. This highlights the ineffectiveness of these attacks in practical VFL scenarios where anomaly detectors are deployed.
@@
(a) Impact of anomaly detector on ASR in VFL inference system. The “Ground-Truth” represents the proportion of the targeted label in the original test set, serving as a baseline for comparison.
@@ 4.2. Main Idea of VTarbel
In summary, we propose a novel two-stage attack framework, VTarbel, which will be described in detail in the next section. This framework generalizes previous methods, where the “Ground-Truth” case and previous attacks illustrated in Figure  4(a) are special cases of our framework. Specifically, the “Ground-Truth” corresponds to using only the first stage without performing attacks in the second stage, while previous attacks correspond to cases where only the second stage is applied without the first stage. In Section  6.3, we demonstrate that both first-stage-only and second-stage-only approaches are ineffective, and only our generalized two-stage framework achieves the highest ASR.
@@ 5.1. Two-stage Attack Formulation
From the above analysis, we identify a fundamental trade-off in the design of the two-stage attack framework. From Eq.(9), it is evident that the total number of test samples predicted as targeted label is the summation of two terms: the number of samples predicted as targeted label in preparation stage and the number predicted in attack stage. If the preparation stage is too short, although a large number of test samples remain for the attack stage, the relatively low ASR $s_{2}$ will result in a small second term in Eq.(9), leading to an ineffective attack. All previous attack methods fall into this category, corresponding to the special edge case where the preparation stage is empty, i.e., $\mathcal{Q}=\emptyset$. On the other hand, if the preparation stage is too long, the ASR $s_{2}$ will be high, but the number of test samples remaining for the attack stage will be small. This also causes the second term in Eq.(9) to be small, resulting in an ineffective attack. The “Ground-Truth” case in Figure  4(a) falls into this category, corresponding to the special edge case where $\mathcal{Q}=D_{adv}$.
The experimental results in Figure  5 confirm our analysis. The overall ASR $s$ initially improves as $\rho$ increases from 1% to 5%. However, when $\rho$ becomes relatively large (10% to 90%), there is a corresponding decrease in $s$. To effectively address this trade-off, our design goal is to construct a preparation stage that maximizes $s_{2}$ while keeping its length minimal, thereby preserving as many test samples as possible for the attack stage.
@@ 6.4. Optimality Verification of Preparation Set
We observe that, in most cases, the preparation set selected by VTarbel achieves the optimal ASR. For instance, under the (ResNet18, CIFAR10)configuration, VTarbel achieves the highest ASR of 90.28%, while the early stopping ($15\%$) and late stopping ($15\%$) variants only achieve ASRs of 81.97% and 83.28%, respectively, both of which are substantially lower than VTarbel’s result. The reason behind this performance difference is as follows: For late stopping, the optimal $|\mathcal{Q}^{*}|$ size provides enough test samples for the surrogate model and estimated detector to be sufficiently trained. Adding more samples in the preparation stage reduces the number of samples predicted as the targeted label during the attack stage, thus lowering the ASR. In contrast, for early stopping, fewer test samples than VTarbel’s optimal preparation set slightly degrade the performance of the estimated detector, leading to a reduced number of samples predicted as targeted label in the attack stage, which also results in a lower ASR compared to VTarbel.
@@ 6.5. Detector-Agnostic Attack
(a) The impact of the attacker’s feature fraction in a two-participant VFL inference system, where the features allocated to each participant are randomly selected.
@@ 6.6. Scalability Evaluation
In this section, we evaluate VTarbel’s attack scalability in terms of the attacker’s feature fraction and the total number of participants in the VFL inference system. Our experiments are conducted using the TabMNIST and TabFMNIST datasets under two scenarios: the first scenario corresponds to a VFL inference system with two participants, where the attacker’s feature fraction (i.e., the fraction of the attacker’s features to the total number of features) ranges from 90% to 10% (Figure 8(a)). The second scenario corresponds to a VFL inference system with varying numbers of participants, ranging from 2 to 10, where the total features are randomly and evenly allocated to each participant (Figure 8(b)).
From both Figure 8(a) and Figure 8(b), several similar trends can be observed. First, the baseline ASR remains stable and is unaffected by the attacker’s feature fraction or the number of participants. This is because the baseline is only related to the global VFL model’s accuracy. Next, as the number of features allocated to the attacker decreases (i.e., the decrease in the attacker’s feature fraction in Figure 8(a) and the increase in the total number of participants in Figure 8(b)), both the accuracy of the locally trained surrogate model and the F1 score of the estimated detector gradually decrease. The reason for this is that with fewer features, it becomes much more challenging for the attacker to leverage these limited features to train an accurate surrogate model or detector that can precisely estimate the global VFL model or the defender’s detector. Consequently, there is a corresponding gradual decrease in VTarbel’s ASR, which is influenced not only by the less accurate surrogate model and detector but also by the benign features from other participants that can neutralize the malicious properties of the attacker’s feature embedding. However, we note that practical VFL systems typically involve only two participants, where the passive party (i.e., the attacker) serves as the feature provider and often possesses more features than the active party (i.e., the defender) (Chen et al., 2021). In such practical settings, VTarbel demonstrates strong attack performance. For instance, as shown in Figure  8(a), when the attacker holds 60% of the features, the ASR reaches 91.97% on TabMNIST and 90.73% on TabFMNIST, posing a significant threat to the security of practical VFL inference systems.
(b) The impact of the number of participants, where the total features are randomly and evenly split among participants.
(c) Impact of an imbalance binary dataset, where the negative class is the targeted label.
@@ 6.7.1. Imbalanced Dataset
We investigate the impact of an imbalanced dataset on VTarbel’s attack performance. We construct a binary test set containing 1,000 samples by randomly selecting test samples from class 0 (treated as the negative class) and class 1 (treated as the positive class) from the CIFAR10 dataset, and evaluate it using the VGG13 and ResNet18 models. As shown in Figure 8(c), the imbalance ratio on the x-axis is defined as the number of negative samples divided by the number of positive samples, where the targeted label is the negative class. For VTarbel, we observe that even as the imbalance ratio increases from 1:1 to 1:8, the ASR remains stable and comparable to that of the balanced dataset. For example, when the imbalance ratio is 1:8, the ASR for VGG13 and ResNet18 is 0.916 and 0.919, respectively, both of which are close to the ASRs of 0.942 and 0.934 observed when the imbalance ratio is 1:1. This demonstrates that VTarbel is robust to class imbalance within the test set. Furthermore, when compared to the variant Random-Preparation (w/o clustering), we find that when the class imbalance becomes more severe (e.g., the imbalance ratio is greater than 1:5), the ASR for Random-Preparation (w/o clustering) is close to zero and even lower than that of the baseline. In contrast, VTarbel consistently achieves a considerably higher ASR than Random-Preparation (w/o clustering). This further validates that the clustering step in the preparation stage is crucial, as it ensures that the preparation set remains balanced, which facilitates the training of both the surrogate model and the estimated detector, even when the original test set is imbalanced.
(d) Impact of anomaly detection threshold under the (ResNet18, CIFAR10, DeepAE) configuration.
@@ 6.7.2. Anomaly Detection Threshold
As shown in Figure 8(d), we plot the overall ASR of VTarbel, along with the ASR of the attack stage (since the detection thresholds have no effect on the ASR of the preparation stage) and the overall anomaly ratio identified by the defender’s detector. In Figure LABEL:fig:active_percentile, as we gradually increase the strength of the defender’s anomaly detector (e.g., with detection threshold percentiles lower than 50), there is a gradual increase in the anomaly ratio. This occurs because most of the attacker’s optimized feature embeddings are recognized as anomalies. As a result, the ASR of the attack stage gradually decreases, and the overall ASR of VTarbel also decreases simultaneously. In contrast, Figure LABEL:fig:passive_percentile shows that as the detection strength of the attacker’s estimated detector increases, the anomaly ratio remains stable and close to 0. The reason for this is that as the detection strength of the estimated detector increases, the optimization step in Algorithm 2 is terminated earlier with fewer optimization rounds, ensuring that the generated feature embedding stays within a region with high probability. Thus, the anomaly ratios consistently remain close to zero. At the same time, due to fewer optimization rounds, the malicious feature embeddings are less effective at being mapped to the targeted label, which explains the slight decrease in the ASR of the attack stage and the overall ASR of VTarbel.